The doctor orders 1/400 grain of Synthroid daily the stock supply is 150 mcg/scored tablet how many tablets will the patient receive?  

A. 10 tabs

B.1.0 tabs

C.10tabs

D.1 tab

E.1/2 tab
Answer #1

ANSWER.

* Calculation of number of tablets will the patient receive.

        Physicians order = 1/400 grain of synthyroid.

            Convert grain to mcg for calculation

                   1 grain = 64798.9 mcg.

        So

            1/400 grain = 161.9 mcg.

          Stock supply = 150 mcg / scored tablet.

   Number of tablet required = 161.9 / 150

                                            = 1.07

                                           D = 1 tab.